OpenStudy (anonymous):

Use signed numbers to solve. A helicopter descended 1200 feet, rose 800 feet, and then descended 450 feet. What was the net gain or loss in altitude? A. lost 2450 ft B. lost 850 ft C. gained 850 ft D. gained 2450 ft

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

A helicopter descended 1200 feet, rose 800 feet, and then descended 450 feet. descended 1200 feet : -1200 rose 800 feet : +800 descended 450 feet : -450 ----------------------------- Overall : ???

ganeshie8 (ganeshie8):

Add them up

OpenStudy (anonymous):

-850

ganeshie8 (ganeshie8):

yes ! that means overall it has "descended" 850 feet.

ganeshie8 (ganeshie8):

which means it has lost a height of 850 feet

ganeshie8 (ganeshie8):

if that makes some sense...

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

can u help me with some more

ganeshie8 (ganeshie8):

shoot...

OpenStudy (anonymous):

Write an equation and then solve the problem. Vick’s Vitamins had start-up expenses of $85,000 for research and $125,000 for the first year’s inventory. So far, the sales are $45,000. What do the sales have to be for the rest of the year for the company to break even? A. $65,000 B. $130,000 C. $165,000 D. $255,000

ganeshie8 (ganeshie8):

at breakeven, there will not be any profit or loss

ganeshie8 (ganeshie8):

expenses = revenues

ganeshie8 (ganeshie8):

85000 + 125000 = 45000+x

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

but 85000 + 125000 dose not = 45000

ganeshie8 (ganeshie8):

good observation :) yes they're not equal

ganeshie8 (ganeshie8):

but did u miss the "+x" attached to the equation, in the end ?

ganeshie8 (ganeshie8):

85000 + 125000 = 45000+x

ganeshie8 (ganeshie8):

you need to solve that "x", such that equation balances

ganeshie8 (ganeshie8):

first add left hand side numbers

OpenStudy (anonymous):

idk how to do it can u help

ganeshie8 (ganeshie8):

85000 + 125000 = 45000+x 210000 = 45000 + x

ganeshie8 (ganeshie8):

next subtract 45000 both sides

ganeshie8 (ganeshie8):

85000 + 125000 = 45000+x 210000 = 45000 + x 165000 = x

ganeshie8 (ganeshie8):

165,000 is the sales they need to make still to breakeven

OpenStudy (anonymous):

where u get the 160000

OpenStudy (anonymous):

165,000 i meant

ganeshie8 (ganeshie8):

did u get my second equaiton ?

ganeshie8 (ganeshie8):

85000 + 125000 = 45000+x 210000 = 45000 + x *********************** 165000 = x

ganeshie8 (ganeshie8):

^^that one, did u understand how I got 210000 ?

OpenStudy (anonymous):

yeah by adding 85000 + 125000

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

To get to third equation, you need to subtract 45000 both sides

ganeshie8 (ganeshie8):

85000 + 125000 = 45000+x 210000 = 45000 + x *********************** -45000 -45000

OpenStudy (anonymous):

u have to add them those -45000 -45000

OpenStudy (anonymous):

@ganeshie8 sorry i am giving u a hard time but i have to do it

ganeshie8 (ganeshie8):

yah subtract : 210000 - 45000 wat do u get ?

OpenStudy (anonymous):

oh 165000

OpenStudy (anonymous):

so the answer is 165000

OpenStudy (anonymous):

The first and second year’s sales for a company were $225,000 and $547,000. The expenses for the first year were $186,000. The company had a net profit of $202,000 after two years. What were the company's expenses for the second year? A. $16,000 B. $384,000 C. $586,000 D. $788,000

OpenStudy (anonymous):

@ganeshie8 sorry but have one more left

OpenStudy (anonymous):

The first and second year’s sales for a company were $225,000 and $547,000. The expenses for the first year were $186,000. The company had a net profit of $202,000 after two years. What were the company's expenses for the second year? A. $16,000 B. $384,000 C. $586,000 D. $788,000

ganeshie8 (ganeshie8):

Profits = Sales - Expenses

ganeshie8 (ganeshie8):

you knw Profits = 202,000

ganeshie8 (ganeshie8):

Can u find out Sales ?

OpenStudy (anonymous):

$225,000 and $547,000

ganeshie8 (ganeshie8):

yes, add them up. wat do u get ?

OpenStudy (anonymous):

$225,000 + $547,000 + 202,000 -186,000

OpenStudy (anonymous):

974000

ganeshie8 (ganeshie8):

Nope, add only the sales first

ganeshie8 (ganeshie8):

Total Sales = $225,000 + $547,000 = ?

OpenStudy (anonymous):

772000 i thought we had to add the profit and sail sorry

ganeshie8 (ganeshie8):

yes, next lets say second year expenses were "x", then total expenses = 186,000 + x

ganeshie8 (ganeshie8):

plug these in the equation : Profit = Sales - Expenses 202,000 = 772000 - 186000 - x

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

idk need help plez @ganeshie8

ganeshie8 (ganeshie8):

202,000 = 772000 - 186000 - x 202,000 = 586000 - x

ganeshie8 (ganeshie8):

I have just subtracted the right hand side terms

ganeshie8 (ganeshie8):

202,000 = 586000 - x next subtract 586000 both sides, you get : -384,000 = -x multiply -1 both sides 384,000 = x

OpenStudy (anonymous):

thanks